### v1

# Investigating the Impact of Trust in Multi-Human Multi-Robot Task Allocation

Abstract
Trust is essential in human-robot collaboration. Even more so in multi-human multi-robot teams where trust is vital to ensure teaming cohesion in complex operational environments. Yet, at the moment, trust is rarely considered a factor during task allocation and reallocation in algorithms used in multi-human, multi-robot collaboration contexts. Prior work on trust in single-human-robot interaction has identified that including trust as a parameter in human-robot interaction significantly improves both performance outcomes and human experience with robotic systems. However, very little research has explored the impact of trust in multi-human multi-robot collaboration, specifically in the context of task allocation. In this paper, we introduce a new trust model, the Expectation Comparison Trust (ECT) model, and employ it with three trust models from prior work and a baseline no-trust model to investigate the impact of trust on task allocation outcomes in multi-human multi-robot collaboration. Our experiment involved different team configurations, including 2 humans, 2 robots, 5 humans, 5 robots, and 10 humans, 10 robots. Results showed that using trust-based models generally led to better task allocation outcomes in larger teams (10 humans and 10 robots) than in smaller teams. We discuss the implications of our findings and provide recommendations for future work on integrating trust as a variable for task allocation in multi-human, multi-robot collaboration.
I Introduction
Trust is crucial in several human-robot collaboration (HRC) tasks, such as search and rescue operations [1, 2], disaster response [3, 4], industrial collaboration [5, 6, 7, 8], and other contexts requiring multi-agent robot collaboration [9, 10, 11, 12, 13, 14, 15]. To capture the different dimensions of trust in HRC, researchers have provided different definitions of trust, including characterizing it as the belief that a human or robot team member will perform as expected in a given task scenario [16]. The guarantee that an agent possesses enough capacity to complete a task in uncertain and dynamic environments [17]. The expectation that agents collaborating on a task or project will not harm each other [18]. Other researchers have also characterized it as the belief that collaborating agents will be willing to act on the information and recommendations given to them to enable them to complete a task [16]. In this research, we align with the definition that posits that trust is the expectation that a team member has the required and expected capacity to complete a task.
Figure 1: A conceptual illustration of the proposed Expectation Comparison Trust (ECT) Model.
Prior work by researchers has adopted different computational strategies for modeling trust, including using probabilistic, non-probabilistic, and graph-based approaches to capture and model the different dynamic and behavioral aspects of trust. However, although several research studies have explored trust in single-HRC, very limited work has explored trust in the context of multi-human multi-robot (MH-MR) collaboration [19, 20, 21, 22]. Considering the increasing proliferation of larger, more diverse MH-MR teams and collaborative activities, it is crucial to investigate how different trust models perform in complex, multi-agent task allocation scenarios and their value in enhancing task allocation outcomes in MH-MR collaboration that involve complex and sensitive tasks.
In this research, we investigated the impact of trust on MH-MR task allocation performance outcomes. Our investigation involved five trust models, including a no-trust baseline model, three models from prior work, Monir et al.[23], the Xu and Dudek Model[24], the Guo and Yang Model [25], and the Expectation Confirmation Trust (ECT) Model we introduce in this research. Our ECT Model draws inspiration from the expectation confirmation theory [26] and applies its core concept for modeling trust in MH-MR contexts. Specifically, the ECT model, as illustrated in Fig. 1, posits that trust is the interplay between the level of divergence in prior expectations of performance and actual performance results, focusing especially on human expectations of robot performance. By conducting this investigation, we seek to highlight the importance of trust in MH-MR contexts and provide insights into the strengths and limitations of different trust modeling approaches in MH-MR contexts. Our research question included: How do different trust models affect task allocation in teams of varying size and composition like 2 humans and 2 robots (2H-2R), 5 humans and 5 robots (5H-5R), and 10 humans and 10 robots (10H-10R)?
Through our research, we make the following contributions: 1) We propose the ECT model to provide robotics researchers and practitioners with the ability to model trust in MH-MR task allocation; 2) We conduct an experiment to investigate the impact of introducing trust in MH-MR teams and found that trust improved the performance of higher team configurations; and 3) We further show that there is no monolithic trust model, as other models performed well in the experiment. This finding shows that there is no monolithic trust model, hence the model suitable for a project depends on various factors, including the task objectives, team composition, and other related factors.
The rest of the paper is structured as follows: Section II provides a detailed background on the evolution of trust models in HRI. Section III describes our proposed trust methodology, the simulation environment, and the case study setup of our experimental design. Section IV presents the results of our comparative analysis, followed by a discussion of the findings and their implications in Section V. Finally, Section VI concludes the paper and outlines directions for future research.
II Related Works
II-A Evolution of Trust Models in Human-Robot Collaboration
Over the past decade, several researchers have explored different approaches for introducing computational approximates of trust into HRC systems, interfaces, and algorithms. The overall goal of this research area is to computationally capture and model the different aspects of human attitudes, behaviors, and heuristics related to trust in ways that can be used to improve performance and safety in HRC. The dimensions of prior work in this area have explored a mixture of techniques for capturing, quantifying, or predicting different trust-approximate variables.
Guo et al. [27] highlighted that trust-enabled human-robot teaming often relies on using trust models to operationalize and transform different ideals of trust into computational approaches that support human-robot teaming. They introduced the reverse psychology trust-behavior model and compared its performance against prior existing models to investigate its performance. Their work further proposed a trust-seeking reward function that balances team performance and trust preferences. Chen et al. [28] proposed trust-POMDP as a means of introducing trust variables into human-robot interaction in ways that improve performance. They tested the performance of their approach using simulation and real-life experiments and found that their approach of factoring in trust improves human-robot team performance. Yemini et al. [29] introduced a mathematical framework for determining teaming consensus, especially where there are malicious agents. They used stochastic trust values for probabilistic measures of trustworthiness between agents in a distributed consensus system. Findings from their research revealed that trust-based approaches played a significant role in enhancing and improving the security and reliability of distributed systems.
Other researchers have also explored trust using different models, including trust-aware motion planning for HRC systems [30], examining the impact of using the Theory of Mind for modeling trust-aware robot policy for HRC [31], investigating the conditions under which robots can trust human wayfinding information for navigational guidance [32], examining approaches for eliciting trust levels during live interaction in HRC [33], and highlighting areas where trust plays a crucial role in robot proficiency by foregrounding to robots the areas where their capabilities are better matched with a given task and vice versa [34]. Relevant here is that although several prior works have explored trust in HRC, very few research have explored trust in the context of complex and sometimes heterogenous MH-MR teams.
II-B Factors Influencing Trust Dynamics
Trust models often seek to quantify the complex interplay of psychological, social, and environmental factors present in HRC. However, researchers have an implicit understanding of the difficulty of achieving a complete model of trust that captures all the rich variables at play in HRC. As a result, prior work by researchers has focused on different trust elements that are peculiar and at play in trust formation and maintenance in a given context of HRC.
Hudspeth et al. [35] conducted a study to investigate the impact of interfaces on human-robot trust collaboration in physical environments. Findings from their study revealed that interfaces and feedback provided through them significantly impact levels of trust in HRC. Luebbers et al. [36] highlighted the role of human memory recall of a robot’s previous performance on trust in HRC, highlighting that a last memory of a robot failing a task leads to increased distrust of humans on robot competence. They conducted an experiment with 53 participants to test this assertion further. Findings from their study validated concerns about the impact of recency bias on human trust in robot competence. Hald and Rehm [37] used several questionnaires, including human-robot trust questionnaires, to elicit information about ways of improving the usability of HRC systems. Findings from their research using such trust questionnaires in systems led to increased usability and high trust in system performance. Rehm et al. [38] highlighted the ways in which trust plays a role in human-robot interaction in social contexts, especially with respect to users relying on or rejecting the assistance of robot-assisted shopping. Findings from their research showed that positive rejection increases user trust in the system.
Furthermore, Wang et al. [39] investigated the value of using trust for role arbitration in situations where control is alternated between humans and robots in ways that optimize team performance. They introduced a quantitative approach for modeling trust and self-confidence in HRC. Findings from their research revealed that their approach delivered superior results in HRC in achieving a task. Xu et al. [40] also investigated the approaches for enhancing trust recognition in HRC using precise approaches. They introduced an EEG-based approach that includes an EEG Vision Transformer and EEGTrust dataset for detecting and measuring levels of human trust during HRC. Findings from their research revealed that their approach produced significant results that outscored other baseline approaches. They further made their code and dataset available to facilitate this genre of research.
III Methods
III-A Trust Models
We employed different trust models for our experiment, including the no-trust baseline model, Monir et al. [23], Xu and Dudek Model [24], Guo and Yang Model [25], and the Expectation Confirmation Model (ECM), which we contribute in through this research. Overall, our objective was to examine the impact of the different trust models on task allocation in the context of MH-MR. We briefly describe each model below.
III-A1 No-trust baseline model
We employed the no-trust baseline model to investigate the outcome of not including trust as a parameter for task allocation in MH-MR contexts and to compare its performance to other trust-enabled models.
III-A2 Monir Trust Model
Monir et al. [23] treated trust as a step function that relies on the performance of the human and robot performance to measure trust levels. They categorized trust into three regions, including unpredictable, predictable, and faithful regions. They highlighted that the human-robot interaction starts with the robot in the dependable region and should ideally end at the faithful region if the interaction goes well. They provided the following equation as a means of modeling their trust concept:
$T(t)=\begin{cases}0,&\text{for }R_{p}(t)<f_{P}\\ \epsilon,&\text{for }f_{P}\leq R_{p}(t)<f_{D}\\ \min(1,\epsilon+\tanh(c\Delta P)),&\text{for }f_{D}\leq R_{p}(t)<f_{F}\\ 1,&\text{for }R_{p}(t)\geq f_{F}\end{cases}$ (1)
where $T$ is the trust level, $t$ is time, $R_{p}(t)$ is the robot’s performance at time $t$, $f_{P}$, $f_{D}$, and $f_{F}$ are thresholds for predictability, dependability, and faith, respectively, $c$ and $\epsilon$ are constants based on the person’s preferences, and $\Delta P=R_{p}(t)-f_{D}$.
III-A3 Xu and Dudek Trust Model
The Xu and Dudek Model [24] used a Dynamic Bayesian Network to model and infer trust in HRC. The model captures different human-robot interaction elements, including human interventions, robot performance, and the overall trust change reports during task completion. The core idea of their model is captured as follows:
$\begin{split}P(t_{k},t_{k-1},p_{k},p_{k-1})&\approx N(t_{k};t_{k-1}+\omega_{tb}+\omega_{tp}p_{k}\\ &\quad+\omega_{td}(p_{k}-p_{k-1}),\sigma_{t})\end{split}$ (2)
$O(t_{k}=1,t_{k-1},i_{k},e_{k}):=S(\omega_{ib}+\omega_{it}t_{k}+\omega_{id}1t_{k}+\omega_{ie}e_{k})$ (3)
$O_{c}(t_{k},t_{k-1},c_{k}):=\text{Prob}(c_{k}|t_{k},t_{k-1})$ (4)
where $N$ denotes a Gaussian distribution, $S(x)$ is the sigmoid function, $\omega_{tb}$, $\omega_{tp}$, and $\omega_{td}$ are weights related to trust, $p_{k}$ is the robot’s performance, $t_{k}$ is the trust state at time $k$, and $i_{k}$ and $e_{k}$ denote human intervention and task change, respectively.
III-A4 Guo and Yang Trust Model
The Guo and Yang model [25] is a personalized trust prediction model that uses a Beta distribution and Bayesian inference to measure trust in HRC. Their model appreciates that trust is not static but evolves as the interaction between the human and robot progresses. They validated their model by comparing it to other trust models that came before it. The core formulation of their model is as follows:
$\alpha_{i}=\begin{cases}\alpha_{i-1}+w_{s},&\text{if }p_{i}=1\\ \alpha_{i-1},&\text{if }p_{i}=0\end{cases}$ (5)
$\beta_{i}=\begin{cases}\beta_{i-1}+w_{f},&\text{if }p_{i}=1\\ \beta_{i-1},&\text{if }p_{i}=0\end{cases}$ (6)
$\hat{t}_{i}=\frac{\alpha_{i}}{\alpha_{i}+\beta_{i}}$ (7)
where $\alpha_{i}$ and $\beta_{i}$ are parameters of the Beta distribution, $p_{i}$ is the performance indicator (1 for success, 0 for failure), $w_{s}$ and $w_{f}$ are weights for positive and negative experiences, and $\hat{t}_{i}$ is the predicted trust at the completion of the $i$-th task.
III-A5 Expectation Confirmation Trust (ECT) Model
Figure 2: Illustration of the core process of our ECT Model.
The ECT Model is built from the Expectation Confirmation Theory, a technique from psychology which defines satisfaction as the interplay between expectation, performance, and disconfirmation [26]. In our context, the ECT model is built on the assumption that trust in the context of HRC is based on the difference between expectations of robot performance before a task and the actual performance of a robot following the completion of a task.
The ECT model addresses a critical gap by introducing trust models for MH-MR teams and explicitly incorporating the dynamic nature of human expectations in collaborative scenarios. Unlike previous single-human robot trust models that often treat trust as a static or slowly evolving parameter, the ECT model recognizes that human trust in robotic teammates is continuously shaped by the interplay between prior expectations and observed performance. This viewpoint is crucial in complex, high-stakes environments where team compositions and task demands change rapidly. By grounding the model in Expectation Confirmation Theory [26], an established framework from consumer psychology, we leverage insights from human decision-making processes that have been extensively studied and adapted in different contexts [41, 42]. This interdisciplinary approach allows us to capture nuanced aspects of trust formation and maintenance in new ways.
The ECT model includes several components as illustrated in Fig. 2, including initial trust, which is often based on the assumption of a robot’s capability and perception; performance evaluation, which is based on the actual performance of a robot on a task; trust updating mechanisms that seek to match initial expectations to performance, and team-trust aggregation function that seeks to determine the overall team trust on a robot capability as a means of measuring and determining the final trust score for task allocation. Our model outline is presented in detail in Table I.
TABLE I: Framework for ECT model for multi-human, multi-robot (MH-MR) systems. This framework addresses human trust in robot performance during MH-MR collaboration.
             Concept                       Definition                       Formula
[…]
Performance Comparison The robot’s actual performance is compared to the expected performance to determine the degree of expectation confirmation. $\Delta_{ir}=P_{ir}-E_{i-1r}$
Interpretation: $\Delta_{ir}>0$: Positive confirmation (robot exceeds expectations). $\Delta_{ir}\leq 0$: Negative or neutral confirmation.
Trust Update Function Trust in robot $r$ is updated based on the performance comparison ($\Delta_{ir}$) using a non-linear function to avoid drastic trust shifts due to isolated events. $T_{ir}=T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\alpha$ is the learning rate, determining how quickly trust adjusts, $\beta$ is the scaling factor that controls the influence of $\Delta_{ir}$ on trust change, and the hyperbolic tangent function ($\tanh$) ensures changes in trust are gradual and bounded.
Trust Decay Trust diminishes over time in the absence of interactions, accounting for the fading impact of past experiences. $T_{ir}=(1-\gamma)\times T_{i-1r}+\alpha\times\tanh(\beta\times\Delta_{ir})$
where $\gamma$ is the decay rate, indicating how quickly trust diminishes without recent interactions, and $T_{ir}=\min(\max(T_{ir},0),1).$
Trust Facets Trust is multi-dimensional, incorporating aspects like competence, ability, and dependability. $T_{ir}=w_{1}T_{\text{com},ir}+w_{2}T_{\text{abi},ir}+w_{3}T_{\text{dep},ir}$
where $w_{1},w_{2}$, and $w_{3}$ are weights reflecting the relative importance of each trust component.
[…]
Allocation Criteria: High Trust: Robots with higher trust levels receive more complex tasks. Low Trust: Robots with lower trust are assigned simpler tasks or monitored more closely.
Team Trust Aggregation Aggregates individual trust assessments to derive a collective trust measure for each robot. $T_{\text{team}r}(t)=\frac{1}{m}\sum_{i=1}^{m}T_{ir}(t)$
which provides an overall trust level that reflects the team’s collective confidence in each robot.
[…]
III-B Simulation
We implemented a simulation environment to evaluate the performance of various trust models in MH-MR task allocation. The simulation employed RL techniques to model agent behavior and decision-making processes. The simulation environment is represented by a 2D grid world of size 100 $\times$ 100 meters. A total of 10 Points of Interest (POIs) are randomly distributed across the map for each simulation run. The environment is defined as:
$E={G,P,T}$ (8)
where $G$ represents the grid world, $P$ is the set of POIs, and $T$ denotes the terrain type. Each POI $p_{i}\in P$ is associated with an attribute $a_{i}\in{{}^{\prime}survivor^{\prime},^{\prime}hazard^{\prime},^{\prime}resource^{\prime}}$, randomly assigned during initialization. Two types of agents are modeled: humans ($H$) and robots ($R$). Robots are further categorized into Unmanned Aerial Vehicles (UAVs) and Unmanned Ground Vehicles (UGVs). Each agent type $j$ has specific attributes:
$A_{j}={v_{j},s_{j}}$ (9)
where $v_{j}$ is the speed (in m/s) and $s_{j}$ is the sensing range (in meters).
Humans: $v_{H}=1.0$ m/s, $s_{H}=10$ m; UAVs: $v_{UAV}=2.0$ m/s, $s_{UAV}=20$ m; UGVs: $v_{UGV}=1.5$ m/s, $s_{UGV}=15$ m. These speeds are chosen to reflect realistic movement capabilities in SAR. The simulation progresses in discrete time steps, where each step represents 1 second of real-time. We implemented and compared five trust models: 1) NoTrust: A baseline model with no trust updates, 2) Monir: Based on Monir et al.’s work, using discrete trust regions, 3) XuDudek: Implementing Xu and Dudek’s continuous trust update mechanism, 4) GuoYang: Utilizing Guo and Yang’s Beta distribution-based trust model, and 5) ECT: Our proposed ECT model incorporates task complexity and robot capability. Tasks are allocated using a Q-learning approach [43]. The state space $S$ is defined by the agent’s distance to the nearest POI and its current trust level. The action space $A$ consists of three possible actions: move to POI, perform task, and wait. The Q-learning update rule is:
$Q(s,a)\leftarrow Q(s,a)+\alpha[r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})-Q(s,a)]$ (10)
where $\alpha=0.1$ is the learning rate, $\gamma=0.9$ is the discount factor, and $r$ is the reward. The exploration rate $\epsilon$ is set to 0.1.
We conducted a series of experiments to evaluate the performance of different trust models across various team compositions and environmental conditions, including A) Small team (2 Humans and 2 Robots; 2H-2R) in flat terrain, B) Medium team (5 Humans and 5 Robots; 5H-5R) in rough terrain, and C) Large team (10 Humans and 10 Robots; 10H-10R) in obstacle-dense terrain. The environments types (e.g., flat terrain) represent levels of task complexity. We used two primary metrics to evaluate performance, including:
  • Task Success Rate (TSR): The proportion of successfully completed tasks.
    $TSR=\frac{\text{Number of completed tasks}}{\text{Total number of tasks}}$ (11)
  • Average Completion Time (ACT): The average time taken to complete tasks.
    $ACT=\frac{\sum_{i=1}^{n}t_{i}}{n}$ (12)
where $t_{i}$ is the completion time of task $i$, and $n$ is the total number of completed tasks.
For each combination of trust model and scenario, we conducted 100 independent simulation runs to ensure statistical significance. Each run lasted for a maximum of 500 time steps or until all tasks were completed. We used the trust questionnaire from Yagoda and Gillian [44] as our theoretical framework for defining trust parameters, specifically focusing on measuring function and performance-related trust parameters, which include competence, ability, and dependability. The simulation logged key metrics at each step, including agent positions, trust levels, task completion times, and task success rates. This data allowed for a comparison of the trust models in terms of their impact on team performance and collaboration in a multi-agent SAR environment.
IV Results and Analysis
(a) Success rate
(b) Average of completion time
Figure 4: Results of the Success Rate and Average Completion Time of the Different Trust Models in team compositions. The star symbol indicates the best performance in each result, the highest success rate, and the shortest completion time. Zoom in for better readability.
Results from our evaluation of the five trust models, including the No Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across three team compositions (e.g., 2H-2R, 5H-5R, and 10H-10R), revealed close results in both task success rates and average task completion times. In small team configurations (2H-2R), we observed that all trust models performed comparably in terms of success rates. ECT, NoTrust, and XuDudek all achieved the highest success rate of 0.35 ± 0.08, while Monir and GuoYang followed closely at 0.33 ± 0.08. Regarding completion times, XuDudek was the fastest at 49.28 ± 16.72 seconds, with ECT following closely at 49.66 ± 14.43 seconds. Notably, ECT demonstrated the lowest standard deviation in completion time, indicating more consistent performance even in small-scale operations. As we moved to medium-sized teams (5H-5R), the performance differences between trust models became more pronounced. Monir took the lead in success rate at 0.63 ± 0.10, with ECT following closely at 0.62 ± 0.10. The other models (NoTrust, XuDudek, and GuoYang) tied at 0.60, albeit with slightly higher standard deviations. Interestingly, the completion time results for this team size revealed a different pattern. Monir was the fastest at 48.42 ± 11.73 seconds, while ECT had the slowest mean completion time at 52.25 ± 11.25 seconds. This suggests a potential trade-off in ECT between maintaining a high success rate and optimizing completion time in moderately complex scenarios.
The most notable results emerged in the large team configuration (10H-10R), where ECT demonstrated its strengths. In this complex scenario, ECT tied with GuoYang for the highest success rate at 0.82 ± 0.10, outperforming the other models. Moreover, ECT achieved the fastest mean completion time of 48.99 ± 9.13 seconds, significantly quicker than the next fastest model, Monir, at 49.14 ± 10.08 seconds. Also, the ECT model exhibited the lowest standard deviations in both success rate and completion time, indicating the most consistent and reliable performance in large-scale MH-MR contexts. An important observation from these results is the scalability demonstrated by the ECT model as its success rate improved from 0.35 in small teams to 0.82 in large teams, representing one of the most significant improvements among all models. This scalability suggests that the ECT model’s trust dynamics become effective as team complexity grows, a critical factor for real-world applications where scalability is often paramount. In the 10H-10R configuration, the ECT model outperformed other models in terms of consistency, demonstrating the lowest standard deviations in both success rate (±0.10) and completion time (±9.13s).
V Discussion
Non-Linearity in Team Size and Performance. Results from our experiment showed a non-linear relationship between team size and performance across the different trust models. This non-linearity aligns with findings from prior work [45] which showed that large autonomous systems often exhibit non-linear behaviors due to the complex interactions between individual robots. The superior performance of trust-based models, particularly the ECT model, in larger team configurations (10H-10R) suggests a potential ’scale threshold’ where trust becomes a critical factor in team efficiency and optimal performance. The findings from our research open up new avenues for research into the optimal team sizing and threshold where trust measurement becomes crucial for MH-MR collaborations.
Trust as a Computational Resource. The ECT’s model’s performance, especially in larger teams introduces a novel perspective of trust as a form of distributed computational resource that leverages the perception of team members vis-a-vis robot performance. Unlike traditional views of trust as a purely psychological construct, our model and the results from the experiment suggest that trust functions as a form of collective intelligence that scales with team size and requires active participation from all team members for an objective outcome. This conceptualization of ’trust as a form of computation’ could help researchers reimagine how we design and implement trust models in robotic systems, potentially leading to an easily scalable and adaptable approach for eliciting and computing trust for large MH-MR teams. The design of the ECT model implicitly means that as the team size increases, the computational power of trust also increases.
Task Allocation and Trust Heuristics in Large Teams. The performance discrepancies between trust-based and non-trust models were more pronounced in larger team configurations and raised important questions about the role of trust in mitigating complexity in MH-MR teaming and task allocation. As team size increases, the computational demands of task allocation grow as well. Our results suggest that trust might serve as an additional heuristic, in addition to other factors like fatigue and workload, to reduce this complexity by providing insight and allowing for more efficient task allocation that is based on empirical metrics in large-scale MH-MR collaboration settings. Specifically, the ECT model is a scalable trust model that is adaptive to different team configurations and sizes and performs relatively well regardless of team composition, showing its usefulness.
Advantages of the ECT Model. Another key benefit of the ECT model lies in its direct approach. While other trust models often rely on independent techniques like facial detection, behavioral cues, and other indirect and after the fact approaches, the ECT model relies on direct elicitation of trust scores from human team members for a task and continuously throughout the lifecycle of the task. This approach ensures that the trust levels of human teams on robot performance are continuously updated. Furthermore, the approach of team trust aggregation allows for congregating the trust scores of all team members in measuring trust levels. Through this approach, the trust rating of an individual team does not have an outsized influence on the overall trust final score generated by the team.
Study Limitations and Future Directions. While results from our study provided valuable insights on trust-based task allocation in MH-MR teams, it is important to also acknowledge the limitations of our study. First, we used simulated environments for our experiment. While this allows for controlled experiments that do not expose users to harm, it may not fully capture the complexities that are often present in real-world contexts. Furthermore, our task domain might not translate to other scenarios and domains. As per prior work, trust is not easy to model, and as a result, our direct question-based method might not capture all the relevant nuances required for measuring trust. Despite these limitations, our work provides a good foundation for future research in trust-based task allocation approaches for MH-MR teams.
VI Conclusion
In this paper, we investigated the impact of trust on task allocation in multi-human multi-robot teams. We employed five trust models, including N-Trust Model, Monir Trust Model, Xu and Dudek Trust Model, Guo and Yang Trust Model, and the ECT Model across different types of teams that included 2H-2R, 5H-5R, and 10H-10R. Our simulation task involved an SAR activity. Findings from our research revealed that trust-based models outperformed the no-trust model most of the time in the 10H10R team configuration and that there was no noticeable difference in the lower configurations. These findings have implications for trust in the context of MH-MR task allocation. Future work could explore trust repair mechanisms for MH-MR teams where trust has been broken due to poor task performance.